# MSP-LLM: A Unified Large Language Model Framework for Complete Material Synthesis Planning

### v2

@@ Abstract
Material synthesis planning (MSP) remains a fundamental and underexplored bottleneck in AI-driven materials discovery, as it requires not only identifying suitable precursor materials but also designing coherent sequences of synthesis operations to realize a target material. Although several AI-based approaches have been proposed to address isolated subtasks of MSP, a unified methodology for solving the entire MSP task has yet to be established. We propose MSP-LLM, a unified LLM-based framework that formulates MSP as a structured process composed of two constituent subproblems: precursor prediction (PP) and synthesis operation prediction (SOP). Our approach introduces a discrete material class as an intermediate decision variable that organizes both tasks into a chemically consistent decision chain. For OP, we further incorporate hierarchical precursor types as synthesis-relevant inductive biases and employ an explicit conditioning strategy that preserves precursor-related information in the autoregressive decoding state. Extensive experiments show that MSP-LLM consistently outperforms existing methods on both PP and SOP, as well as on the complete MSP task, demonstrating an effective and scalable framework for MSP that can accelerate real-world materials discovery.
@@ 1 Introduction
In real-world applications, MSP is explored within a materials space defined at the level of chemical compositions (Na, 2023; Okabe et al., 2024; Choi et al., 2025), and requires predicting a set of precursor materials PP and a sequence of synthesis operations $O$ for a given desired product material. Here, $P\in 2^{\mathcal{M}}\setminus\{\emptyset\}$ denotes a set of precursor materials, where $\mathcal{M}$ denotes the set of materials, and $O\in\mathcal{O}^{*}$ denotes a finite-length synthesis operation sequence. Formally, the MSP task can be defined as $f:\mathcal{M}\rightarrow 2^{\mathcal{M}}\times\mathcal{O}^{*}$, which maps a given product material to a set of precursor materials and a corresponding synthesis operation sequence. However, due to the inherent challenges of MSP, most existing studies have addressed only partial aspects of MSP, rather than providing a framework that solves MSP in its entirety. In particular, prior works typically focus on one of the following two subproblems: (1) precursor prediction (PP), which predicts a set of precursor materials from a given product material, $f_{p}:\mathcal{M}\rightarrow 2^{\mathcal{M}}$, or (2) synthesis operation prediction (OP), which aims to predict a sequence of synthesis operations given a product material and a set of precursor materials, $f_{o}:\mathcal{M}\times 2^{\mathcal{M}}\rightarrow\mathcal{O}^{*}$.
Among these subproblems, OP remains largely underexplored and has been studied only in narrowly scoped settings, typically restricted to specific material systems rather than general MSP (Na, 2023; Pan et al., 2025). By contrast, most existing efforts have focused on PP, yet these approaches often rely on complex retrieval pipelines (He et al., 2023; Noh et al., 2024) or access to proprietary LLM APIs (Kim et al., 2024; Prein et al., 2025b), which limits their scalability and reproducibility. These limitations underscore the need for the scalable and unified modeling framework that can address both OP and PP within a general MSP setting.
@@
However, despite their promise, effective strategies for fine-tuning LLMs to address MSP remain largely unexplored. Specifically, no existing work has established a principled fine-tuning strategy that enables LLMs to perform PP and OP in a unified MSP framework.
In this paper, we propose a unified LLM-based framework, called MSP-LLM, for solving complex MSP tasks. We develop novel task-specific LLM fine-tuning strategies for the two constituent subproblems of MSP, i.e., PP and OP. Specifically, (i) we introduce a discrete material class variable that captures synthesis-relevant characteristics of materials and employ it in both PP and OP as an intermediate decision variable. By incorporating this variable, each subtask is organized into a structured decision chain, in which the model conditions its task-specific generation on the material class variable, leading to more accurate predictions. Furthermore, for SOP, we introduce an explicit conditioning strategy that integrates hierarchical precursor types as synthesis-relevant inductive biases and enforces their effective utilization through precursor constraint factorization, promoting precursor-conditioned operation generation and more reliable retention of precursor constraints in the decoder state. By integrating LLMs fine-tuned for each subproblem within a single coherent framework, MSP-LLM enables complete MSP from target material compositions. Extensive experiments demonstrate that MSP-LLM significantly outperforms existing methods on both individual subproblems and the overall MSP task, highlighting its effectiveness and practical applicability to real-world materials discovery.
@@ 5 Experiments
We evaluate the effectiveness of MSP-LLM  across three settings: the PP task (Section 5.1), which predicts precursor sets from a target material; the SOP task (Section 5.2), which predicts synthesis operations given the target material and ground-truth precursors; and the complete MSP task (Section 5.3), which predicts both precursors and synthesis operations simultaneously from the target material alone.

### v3

# arXiv:2602.07543v3[cs.AI] 01 Mar 2026

@@ Abstract
Material synthesis planning (MSP) remains a fundamental and underexplored bottleneck in AI-driven materials discovery, as it requires not only identifying suitable precursor materials but also designing coherent sequences of synthesis operations to realize a target material. Although several AI-based approaches have been proposed to address isolated subtasks of MSP, a unified methodology for solving the entire MSP task has yet to be established. We propose MSP-LLM, a unified LLM-based framework that formulates MSP as a structured process composed of two constituent subproblems: precursor prediction (PP) and synthesis operation prediction (SOP). Our approach introduces a discrete material class as an intermediate decision variable that organizes both tasks into a chemically consistent decision chain. For SOP, we further incorporate hierarchical precursor types as synthesis-relevant inductive biases and employ an explicit conditioning strategy that preserves precursor-related information in the autoregressive decoding state. Extensive experiments show that MSP-LLM consistently outperforms existing methods on both PP and SOP, as well as on the complete MSP task, demonstrating an effective and scalable framework for MSP that can accelerate real-world materials discovery.
@@ 1 Introduction
In real-world applications, MSP is explored within a materials space defined at the level of chemical compositions (Na, 2023; Okabe et al., 2024; Choi et al., 2025), and requires predicting a set of precursor materials PP and a sequence of synthesis operations $O$ for a given desired product material. Here, $P\in 2^{\mathcal{M}}\setminus\{\emptyset\}$ denotes a set of precursor materials, where $\mathcal{M}$ denotes the set of materials, and $O\in\mathcal{O}^{*}$ denotes a finite-length synthesis operation sequence. Formally, the MSP task can be defined as $f:\mathcal{M}\rightarrow 2^{\mathcal{M}}\times\mathcal{O}^{*}$, which maps a given product material to a set of precursor materials and a corresponding synthesis operation sequence. However, due to the inherent challenges of MSP, most existing studies have addressed only partial aspects of MSP, rather than providing a framework that solves MSP in its entirety. In particular, prior works typically focus on one of the following two subproblems: (1) precursor prediction (PP), which predicts a set of precursor materials from a given product material, $f_{p}:\mathcal{M}\rightarrow 2^{\mathcal{M}}$, or (2) synthesis operation prediction (SOP), which aims to predict a sequence of synthesis operations given a product material and a set of precursor materials, $f_{o}:\mathcal{M}\times 2^{\mathcal{M}}\rightarrow\mathcal{O}^{*}$.
Among these subproblems, SOP remains largely underexplored and has been studied only in narrowly scoped settings, typically restricted to specific material systems rather than general MSP (Na, 2023; Pan et al., 2025). By contrast, most existing efforts have focused on PP, yet these approaches often rely on complex retrieval pipelines (He et al., 2023; Noh et al., 2024) or access to proprietary LLM APIs (Kim et al., 2024; Prein et al., 2025b), which limits their scalability and reproducibility. These limitations underscore the need for the scalable and unified modeling framework that can address both SOP and PP within a general MSP setting.
@@
However, despite their promise, effective strategies for fine-tuning LLMs to address MSP remain largely unexplored. Specifically, no existing work has established a principled fine-tuning strategy that enables LLMs to perform PP and SOP in a unified MSP framework.
In this paper, we propose a unified LLM-based framework, called MSP-LLM, for solving complex MSP tasks. We develop novel task-specific LLM fine-tuning strategies for the two constituent subproblems of MSP, i.e., PP and SOP. Specifically, (i) we introduce a discrete material class variable that captures synthesis-relevant characteristics of materials and employ it in both PP and SOP as an intermediate decision variable. By incorporating this variable, each subtask is organized into a structured decision chain, in which the model conditions its task-specific generation on the material class variable, leading to more accurate predictions. Furthermore, for SOP, we introduce an explicit conditioning strategy that integrates hierarchical precursor types as synthesis-relevant inductive biases and enforces their effective utilization through precursor constraint factorization, promoting precursor-conditioned operation generation and more reliable retention of precursor constraints in the decoder state. By integrating LLMs fine-tuned for each subproblem within a single coherent framework, MSP-LLM enables complete MSP from target material compositions. Extensive experiments demonstrate that MSP-LLM significantly outperforms existing methods on both individual subproblems and the overall MSP task, highlighting its effectiveness and practical applicability to real-world materials discovery.
@@ 5 Experiments
o We evaluate the effectiveness of MSP-LLM  across three settings: the PP task (Section 5.1), which predicts precursor sets from a target material; the SOP task (Section 5.2), which predicts synthesis operations given the target material and ground-truth precursors; and the complete MSP task (Section 5.3), which predicts both precursors and synthesis operations simultaneously from the target material alone.